40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1188     Entry time: Mon Dec 8 17:50:21 2008
Author: Yoichi 
Type: Update 
Category: SUS 
Subject: ITMY drift 
The suspension drift monitor shows that the ITMY alignment was shifted after the earthquake.
Looks like only the UL sensor had a step at the earthquake (see the attachment 1).
So it is probably an electronics problem.
I pushed in the cable between the rack and the ITMY satellite amplifier, but no change observed.
Actually, the ITMY-UL sensor looks like it has been dead before the earthquake.
The second attachment shows a long-term trend of the UL sensor.
The sensor output had been around zero since Nov. 17th.
When I disabled the output of the UL sensor, the sus-drift-mon fields turned green.
So I think the drift-mon's reference values are wrong, and currently the ITMY is in a good alignment.

I also attached the free-swing measurements of the ITMY taken on Aug. 18th and today.
There is no notable change in the resonant frequencies.
Attachment 1: ITMY-OSEMs.png  75 kB  | Show | Hide all | Show all
Attachment 2: ITMY-UL.png  57 kB  | Show | Hide all | Show all
Attachment 3: ITMY-08-18.pdf  160 kB  | Hide | Hide all | Show all
ITMY-08-18.pdf
Attachment 4: ITMY-12-08.pdf  158 kB  | Hide | Hide all | Show all
ITMY-12-08.pdf
ELOG V3.1.3-